OpenStudy (anonymous):

what is the derivative of A*exp(6x)

OpenStudy (anonymous):

\[6A\exp(6x)\] by the chain rule

OpenStudy (theviper):

\[\Huge{\color{red}{\text{Welcome to OPENSTUDY}\quad\color{blue}{\ddot{\smile}}}}\]

OpenStudy (anonymous):

Hint: \(\dfrac{d}{dx} e^u = e^u \dfrac{du}{dx}\)